OpenStudy (anonymous):

What are the real and extraneous solutions in these 2 equations: 1. (x+5)/(x-3)=x 2. (2x)/(x^2)=(x-6)/(-4)

OpenStudy (anonymous):

1. (x+5) = x^2 - 3x x^2 - 4x - 5 = 0 (x-5)(x+1) so x = 5 and x = -1 neither are extraneous (both can substitute into the original)

OpenStudy (anonymous):

wait what about number 2?

OpenStudy (anonymous):

If you cross multiply -8x = x^3 - 6x^2 0 = x^3 - 6x^2 + 8x x = 0 x^2 - 6x + 8 = 0 (x -4)(x - 2) = 0 x = 0, 2, 4 but x = 0 is extraneous (you cannot substitute it into the original problem without having an issue 0/0).